OpenStudy (anonymous):

Please help?! <3 Solve x – 2y = 6 3x – 6y = 18 (2, –2) (3, negative three halves) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

plan b @blondie16

OpenStudy (anonymous):

haha. ok

OpenStudy (anonymous):

lol reaper xD... solve for y then graph it

zepdrix (zepdrix):

So you can either solve by Elimination or Substitution. It looks like Elimination will work really well for this problem. :D \[\large x-2y=6\]\[\large 3x-6y=18\]If we multiply the first equation by 3, we'll get,\[\large \color{#3366CF}{(3)} (x-2y)=(6)\color{#3366CF}{(3)}\]Which becomes,\[\large 3x-6y=18\]See how the first and second equation are exactly the same now??? :O

OpenStudy (anonymous):

@zepdrix she has to graph it though

OpenStudy (anonymous):